40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2474     Entry time: Tue Dec 10 16:35:03 2019
Author: Aidan 
Type: Laser 
Category: 2micronLasers 
Subject: Calibrated cryo-PD REF PD output in units of power incident on cryo chamber window 

We placed a power meter after the fiber collimator, 75mm focal length lens and HR mirror at 45 degrees - basically, we placed the power meter immediately before the input window to the cryo chamber after all the intervening optics from the fiber output.

For a series of laser diode current levels, we measured the power on the power meter and the corresponding voltage on the reference photodetector that is monitoring a 10% pick-off from the laser. The calibration is as follows:

Diode current (mA) REF PD (V) Laser Power (mW)
100.83 1.048 1.171
90.75 0.941 1.043
80.66 0.827 0.904
70.56 0.708 0.763
60.47 0.596 0.627
50.45 0.48 0.485
39.92 0.357 0.338
29.82 0.237 0.191
19.73 0.117 0.045
9.64 0.080 0.000
0.00 0.080 0.000

 

ELOG V3.1.3-